OpenStudy (anonymous):

If SinX=3/5 on the interval (π/2, π), find the exact value of Tan2X. @wet and @shelbygt520 Last question for the night!

OpenStudy (anonymous):

tbh I have no idea how to do this. this might be nonsense

OpenStudy (anonymous):

that's what i did but i don't know how to "find tan2X"

OpenStudy (anonymous):

What's the 2 do?

OpenStudy (anonymous):

it's either \[\large \tan(2x) \]\[\large \tan^2x \]

OpenStudy (anonymous):

I'm betting on the latter. so find tan x and square it

OpenStudy (anonymous):

sorry boys I gtg I have the night shift tonight ;p ill try not to get shot lol

OpenStudy (anonymous):

a working girl *~* Have fun. good night. sweet dreams. etc.

OpenStudy (anonymous):

awmigawd, teacher said it's tan2x

OpenStudy (anonymous):

there's probs a formula for tan(2x). I really don't know this though, sorry

OpenStudy (anonymous):

sin x = 3/5. x between Pi/2 and Pi. First, find cos x --> cos^2 x = 1 - 9/25 = 16/25 --> cos x = -4/5 sin 2x = 2sin x*cos x = -2(3/5)(4/5) = -...... cos 2x = 1 - 2sin^2 x = 1 - 2(9/25) = ..... tan 2x = sin 2x/cos 2x = -.......